OpenStudy (anonymous):

integration of (e^tanx)/(cos^2)x a=0 b=pie/4

OpenStudy (anonymous):

try doing a substitution, u = tan x.

OpenStudy (anonymous):

∫ [e^(tan x)] / cos ^ 2 (x) dx, from 0 to pi/4 = ∫ [e^(tan x)] d tan x, from 0 to pi/4 = [e^(tan x)], from 0 to pi/4 = e - 1 --------- Ideas: (tan x)' = 1/cos^2 (x) and use mental substitution.